# Compliance-to-Code: Enhancing Financial Compliance Checking via Code Generation

### v1

@@ Abstract
Nowadays, regulatory compliance has become a cornerstone of corporate governance, ensuring adherence to systematic legal frameworks. At its core, financial regulations often comprise highly intricate provisions, layered logical structures, and numerous exceptions, which inevitably result in labor-intensive or comprehension challenges. To mitigate this, recent Regulatory Technology (RegTech) and Large Language Models (LLMs) have gained significant attention in automating the conversion of regulatory text into executable compliance logic. However, their performance remains suboptimal particularly when applied to Chinese-language financial regulations, due to three key limitations: (1) incomplete domain-specific knowledge representation, (2) insufficient hierarchical reasoning capabilities, and (3) failure to maintain temporal and logical coherence. One promising solution is to develop a domain specific and code-oriented datasets for model training. Existing datasets such as LexGLUE, LegalBench, and CODE-ACCORD are often English-focused, domain-mismatched, or lack fine-grained granularity for compliance code generation. To fill these gaps, we present Compliance-to-Code  [^1] https://github.com/AlexJJJChen/Compliance-to-Code, the first large-scale Chinese dataset dedicated to financial regulatory compliance. Covering 1,159 annotated clauses from 361 regulations across ten categories, each clause is modularly structured with four logical elements-subject, condition, constraint, and contextual information-along with regulation relations. We provide deterministic Python code mappings, detailed code reasoning, and code explanations to facilitate automated auditing. To demonstrate utility, we present FinCheck: a pipeline for regulation structuring, code generation, and report generation. Compliance-to-Code establishes a new benchmark for LLM-based compliance automation. Experimental evaluation shows that Qwen3-8B substantially improves both structural parsing and compliance code generation capabilities with Supervised Fine-tuning.
@@ 1 Introduction
Dataset Construction and Features: The Compliance-to-Code dataset is built from a systematic analysis of 361 authoritative Chinese financial regulations, spanning ten regulatory categories. We extract and decompose 1,159 representative legal clauses into four fundamental logical components-subject, condition, constraint, and contextual information-to create fine-grained Compliance Units (CUs). We also annotate relationships among CUs to support advanced compositional and cross-reference reasoning. Crucially, each structured CU is paired with a corresponding executable Python module (307 in total), which interacts with a simulated public company database for compliance auditing. Each code example includes detailed Chain-of-Thought reasoning and explanatory comments, enhancing transparency and reproducibility.
@@
Experimental Results: We benchmark recent LLMs for regulation structuring and code generation tasks using our dataset’s detailed annotations. Supervised fine-tuning shows that Qwen3-8B (THUDM, 2025) excels at regulation structuring and notably improve compliance-checking code generation.
@@ 5.4 Training Protocols
Base Models: For Task 1, we utilize Qwen3-8B (Qwen Team, 2025), DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), GLM-4-9B-0414 (THUDM, 2025), GPT-4.1 (OpenAI, 2025), Gemini-pro 2.5 (Kavukcuoglu, 2025), and Claude-3.7 (Anthropic, 2025). For Tasks 2 and 3, which focus on code generation, we additionally include Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) and DeepSeek-Coder-7B-Instruct (Guo et al., 2024) in our evaluation.
@@ 6.1 Task 1 Performance: Advancing Automated Structuring of Complex Regulatory Text
Our evaluation demonstrates that domain-adapted supervised fine-tuning is indispensable for structuring complex regulatory text into actionable CUs and inter-unit relations. The fine-tuned GLM-4-9B achieves 71.7% F1 in CU field extraction and 78.4% for boundary identification, outperforming strong few-shot baselines such as GPT-4.1 by over 8 absolute F1 points.
@@ 6.2 Tasks 2&3: Structured Regulations Are Transformative for Compliance Code Generation
Few-shot prompting and chain-of-thought (CoT) approaches consistently outperform zero-shot and direct prompting baselines, especially on more challenging examples. For instance, in the difficult subset, CoT-LLM (one-shot) achieves up to 34.09% Pass@1 (Task 2), compared to near-zero for zero-shot methods. This underscores the value of providing exemplars and reasoning scaffolds, particularly for complex or ambiguous compliance scenarios.
@@ 7 Conclusion
We have introduced the Compliance-to-Code Dataset and a novel structured synthesis pipeline, embodied in our FinCheck system, for executable compliance checking in the complex domain of financial regulation. Our approach, which formalizes regulatory logic as annotated Compliance Units and translates these into verifiable Python code, demonstrates significant gains in functional accuracy (achieving up to 44.44% Pass@1 with gold structures, a 2.6 times improvement over CoT prompting on end-to-end tasks), interpretability through traceable commentary, and inherent auditability compared to end-to-end LLM prompting. This research not only provides a valuable new resource and strong benchmark results but also opens new avenues for developing safer, more scalable, and reliable RegTech solutions.
@@ Appendix A Liscense
The regulatory documents published by the Beijing Stock Exchange (BSE), used to create, are publicly published.
@@ Appendix F Annotation Schema: Compliance Units and Inter-Unit Relations
Regulatory provisions are rarely isolated; they interact logically. We annotate Inter-Unit Relations between Compliance Units to capture these dependencies, enabling the construction of a directed graph representing the legal logic flow. The four primary relation types annotated are:
@@ Appendix G Error Analysis
  • Upstream Error Propagation (approx. 45% of errors): The most significant error source is inaccuracies from Task 1. Incorrectly predicted inter-unit relations (e.g., missing an exclude condition, or misclassifying refer to as should include) lead to flawed logical flow in the generated code. Erroneous CU boundary detection can also truncate or conflate conditions.
@@
This detailed error breakdown not only highlights current limitations but also provides a clear roadmap for future research, emphasizing the need for more robust relational reasoning in Task 1 and advanced compositional generalization in the code generation phase for complex multi-CU scenarios.

### v2

@@ Abstract
Nowadays, regulatory compliance has become a cornerstone of corporate governance, ensuring adherence to systematic legal frameworks. At its core, financial regulations often comprise highly intricate provisions, layered logical structures, and numerous exceptions, which inevitably result in labor-intensive or comprehension challenges. To mitigate this, recent Regulatory Technology (RegTech) and Large Language Models (LLMs) have gained significant attention in automating the conversion of regulatory text into executable compliance logic. However, their performance remains suboptimal particularly when applied to Chinese-language financial regulations, due to three key limitations: (1) incomplete domain-specific knowledge representation, (2) insufficient hierarchical reasoning capabilities, and (3) failure to maintain temporal and logical coherence. One promising solution is to develop a domain specific and code-oriented dataset for model training. Existing datasets such as LexGLUE, LegalBench, and CODE-ACCORD are often English-focused, domain-mismatched, or lack fine-grained granularity for compliance code generation. To fill these gaps, we present Compliance-to-Code  [^1] https://github.com/AlexJJJChen/Compliance-to-Code, the first large-scale Chinese dataset dedicated to financial regulatory compliance. Covering 1,159 annotated clauses from 361 regulations across ten categories, each clause is modularly structured with four logical elements-subject, condition, constraint, and contextual information-along with regulation relations. We provide deterministic Python code mappings, detailed code reasoning, and code explanations to facilitate automated auditing. To demonstrate utility, we present FinCheck: a pipeline for regulation structuring, code generation, and report generation. Compliance-to-Code establishes a new benchmark for LLM-based compliance automation. Experimental evaluation shows that GLM-4-9B-0414 achieves the best performance on key task in regulation structuring and Qwen3-8B performs the best in compliance code generation tasks.
@@ 1 Introduction
Dataset Construction and Features: The Compliance-to-Code dataset is built from a systematic analysis of 361 authoritative Chinese financial regulations, spanning ten regulatory categories. We extract and decompose 1,159 representative legal clauses into four fundamental logical components-subject, condition, constraint, and contextual information-to create fine-grained Compliance Units (CUs). We also annotate relationships among CUs to support advanced compositional and cross-reference reasoning. Crucially, we selected CUs amenable to computational assessment, annotated them with executable Python modules, and used these as templates to create 307 coding tasks. Each code example includes detailed Chain-of-Thought reasoning and explanatory comments, enhancing transparency and reproducibility.
@@
Experimental Results: We benchmark recent LLMs for regulation structuring and code generation tasks using our dataset’s detailed annotations. Qwen3-8B (Yang et al., 2025) achieves the best performance on code generation. With Supervised Fine-Tuning (SFT), GLM-4-9B-0414 notably improves its regulation structuring capabilities.
@@ 5.4 Training Protocols
Base Models: For Task 1, we utilize Qwen3-8B (Qwen Team, 2025), DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), GLM-4-9B-0414 (THUDM, 2025), GPT-4.1 (OpenAI, 2025), Gemini-pro 2.5 (Kavukcuoglu, 2025), and Claude-3.7 (Anthropic, 2025). For Tasks 2 and 3, which focus on code generation, we additionally include Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) and DeepSeek-Coder-6.7B-Instruct (Guo et al., 2024) in our evaluation.
@@ 6.1 Task 1 Performance: Advancing Automated Structuring of Complex Regulatory Text
Our evaluation demonstrates that domain-adapted supervised fine-tuning is indispensable for structuring complex regulatory text into actionable CUs and inter-unit relations. The fine-tuned GLM-4-9B-0414 achieves 71.7% F1 in CU field extraction, outperforming strong few-shot baselines such as GPT-4.1 by nearly 8 absolute F1 points.
@@ 6.2 Tasks 2&3: Structured Regulations Are Transformative for Compliance Code Generation
Few-shot prompting and chain-of-thought (CoT) approaches generally outperform zero-shot and direct prompting baselines, especially on more challenging examples. For instance, in the difficult subset, CoT-LLM (one-shot) achieves up to 34.09% Pass@1 (Task 2), compared to near-zero for zero-shot methods. This underscores the value of providing exemplars and reasoning scaffolds, particularly for complex or ambiguous compliance scenarios.
@@ 7 Conclusion
We have introduced the Compliance-to-Code Dataset and a novel structured synthesis pipeline, embodied in our FinCheck system, for executable compliance checking in the complex domain of financial regulation. Our approach, which formalizes regulatory logic as annotated Compliance Units and translates these into verifiable Python code, demonstrates significant gains in functional accuracy (achieving up to 78.1% Pass@1 with gold structures and reasoning steps, interpretability through traceable commentary, and inherent auditability compared to end-to-end LLM prompting. This research not only provides a valuable new resource and strong benchmark results but also opens new avenues for developing safer, more scalable, and reliable RegTech solutions.
@@ Appendix A License
The regulatory documents published by the Beijing Stock Exchange (BSE), used to create the dataset, are publicly available.
@@ Appendix F Annotation Schema: Compliance Units and Inter-Unit Relations
Regulatory provisions are not usually isolated; they interact logically. We annotate Inter-Unit Relations between Compliance Units to capture these dependencies, enabling the construction of a directed graph representing the legal logic flow. The four primary relation types annotated are:
@@ Appendix G Error Analysis
  • Upstream Error Propagation (approx. 45% of errors): The most significant error source is inaccuracies from Task 1. Key issues include errors in partitioning regulatory text into its Condition and Constraint components, or failing to fully capture and propagate essential contextual information. Erroneous CU boundary detection can also truncate or conflate conditions.
@@
This detailed error breakdown not only highlights current limitations but also provides a clear roadmap for future research. It emphasizes the need for more robust Task 1 models and advanced compositional generalization in the code generation phase. To effectively handle complex multi-CU scenarios, developing stronger relational reasoning capabilities will also be essential.